OpenStudy (anonymous):

if a^3 = b^3 + c^3 + d^3 , then least value of a is ?

OpenStudy (anonymous):

PLS HELP

OpenStudy (anonymous):

hi kushashwa,,Are you in algebra class or calculus?

OpenStudy (anonymous):

AALGEBRA

OpenStudy (anonymous):

PLS HELP ME

OpenStudy (anonymous):

HELP ME PLEASE

OpenStudy (anonymous):

a<0 , b<0 , c<0

OpenStudy (anonymous):

HOW

OpenStudy (watchmath):

Assuming that a,b,c,d are all positive then we have \(a^3\geq 3\sqrt[3]{b^3c^3d^3}=3bcd\) Hence \[a\geq \sqrt[3]{3bcd}\] So \(\sqrt[3]{3bcd}\) is the minimum value of \(a\).

OpenStudy (anonymous):

OK THANKS VERY MUCH

OpenStudy (anonymous):

ANY1 HAVE A DIFFERENT ANSWER????????????///

OpenStudy (anonymous):

WATCH MATH UR ANSWER IS NOT IN OPTION A)3 B)9 C)6 D) 12

OpenStudy (watchmath):

Then you didn't put the problem completely. What do we know about a,b,c,d? we need more context.

OpenStudy (anonymous):

no it is the full given

OpenStudy (anonymous):

help me any1 else

OpenStudy (watchmath):

The complete problem usually use a complete sentece like this: Let \(a,b,c,d\) be such and such .....

OpenStudy (anonymous):

can we do it with hit and trial method

OpenStudy (anonymous):

i thnk it is , if b=c=d then a^3 = b^3 + b^3 + b^3 =3b^3

OpenStudy (anonymous):

i think we should refer options first

OpenStudy (anonymous):

now if b=1, then a^3=3(1)^3 a^3=3

OpenStudy (anonymous):

hey the answer is 9 i think

OpenStudy (anonymous):

because by hit and trial method 9 is the least number that satisfies the given condition 9^3 = 8^3 + 6^3 + 1^3 729 = 512 + 216 + 1 729 = 729

OpenStudy (anonymous):

WHAT DO U ALL SAY ABOUT THIS ANSWER I THINK I M CORRECT

OpenStudy (anonymous):

PLS SUGGEST

OpenStudy (anonymous):

because by hit and trial method 9 is the least number that satisfies the given condition 9^3 = 8^3 + 6^3 + 1^3 729 = 512 + 216 + 1 729 = 729 IS THIS CORRECT SOLUTION

OpenStudy (anonymous):

pls reply me

OpenStudy (anonymous):

i just did hit and trial method

OpenStudy (anonymous):

but where did you get those 8, 6 and 1?

OpenStudy (anonymous):

if a^3 = b^3 + c^3 + d^3 hence i thought that a must be greater than b, c and d. so if a = 9 hence b , c and d will be among 1, 2, 3, 4 , 5, 6, 7, 8 i took the cubes of these numbers and took the numbers whose sum contains unit digit as 9.

OpenStudy (anonymous):

hence i found many pairs but i found 8, 6 and 1 as correct group . AM I CORRECT PLS TELL ME

OpenStudy (anonymous):

i think it is 3,,, try b=1, c=0 d=0, then try c=1,b=0,d=0, then try d=1,b=0,c=0

OpenStudy (anonymous):

M VERY SORRY TO ALL OF U THE ANSWER WILL BE 6 6^3 = 5^3 + 4^3 + 3^3 216 = 216

OpenStudy (anonymous):

NO MARK IT CAN NOT BE THREE 3^3 IS NOT EQUAL TO ANY SUM OF THREE CUBES

OpenStudy (anonymous):

lol...well thats the least i can see ...lol,, so accordingly b cant be equal to c cant be equal to d,....ok....we need to complete the prob specifically...lol

OpenStudy (anonymous):

TELL ME CLEARLY WHAT IS THE ANSWER. WHAT DO U THINK. I THINK 6 WOULD BE THE ANSWER

OpenStudy (anonymous):

A)3 B)9 C)6 D) 12

OpenStudy (anonymous):

say b=1,c=2 and d=3... also try b-0,c=1 and d=3

myininaya (myininaya):

3 since 3^3=3^3+0^3+0^3

OpenStudy (anonymous):

yes thats it its 3

OpenStudy (anonymous):

YEAH THAT MAY BE

myininaya (myininaya):

it is since there are no restrictions on a,b,c,d, right?

OpenStudy (anonymous):

wait

myininaya (myininaya):

i mean of the options for a that is the smallest but there can be smaller

myininaya (myininaya):

like a->-infinity would be the smallest lol since a negative to an odd power is still negative and can be negative large

OpenStudy (anonymous):

but it is showing 6 as the answer

myininaya (myininaya):

do b,c,d have to be postiive?

myininaya (myininaya):

if not then 6 is wrong

OpenStudy (anonymous):

i also don't know the question that was given i typed i think the question is wrong it should mention the whole context thanks very much

OpenStudy (anonymous):

ok, wc and thnx